# Diagnosing the AGN population origin of TeV neutrinos with their spatial correlation

### v1

@@ IV Predictive Power for Future Neutrino Detectors
Figure 2 illustrates how the detection significance depends jointly on the angular resolution and the total number of detected neutrinos, using a sample of events with reconstructed energies above 100 TeV, where the astrophysical neutrino fraction is fixed at $f=0.5$. The trends are clear: improving the angular resolution reduces positional uncertainties, thereby increasing the contrast between AGN–neutrino associations and random coincidences. At a fixed angular resolution, increasing the total event count enhances the statistical weight of the likelihood, roughly following the expected ${(S/N)}_{f}\propto\sqrt{N}$ scaling. Conversely, when the number of detected neutrinos is fixed, the significance scales approximately with the inverse of the PSF width, i.e. ${(S/N)}_{f}\propto\sigma_{\rm psf}^{-1}$. Notably, achieving a $5\sigma$ detection requires only $\sim 2000$ events when the angular resolution is as fine as $0.1^{\circ}$, whereas resolution coarser than $0.4^{\circ}$ demands orders of magnitude more events, making such significance practically unattainable. At present, IceCube’s effective exposure and angular resolution limit the attainable significance: with $\sim 10^{2}$$10^{3}$ detected astrophysical neutrinos and a median angular uncertainty of $\sim 0.55^{\circ}$ for above 100 TeV (IceCat-2), the current data fall short of the regime where AGN population correlations could be statistically established (denoted by ”IceCube” in Figure 2). This highlights angular resolution as the dominant driver of sensitivity. This trend can be explained as arising from two effects: (i) a narrower $\sigma_{\rm psf}$ allows inclusion of a larger portion of the AGN population—including contributions from fainter sources that would otherwise be lost in confusion noise; and (ii) improved localization concentrates the signal probability density closer to the true source positions, thereby increasing the likelihood ratio and boosting the test statistic.
@@
Figure 3: Trade-off between neutrino energy threshold and exposure time for future neutrino observatories. The left panel shows results for IceCube-Gen2, adopting the projected energy-dependent angular resolution from ice2_psf2019. The right panel shows the corresponding results for the proposed 30 km3 HUNT detector, incorporating its energy-dependent angular resolution and effective area (private communication with Tian-Qi Huang).
@@
Building upon the angular resolution-based considerations described above, we next examine how the selection of different neutrino energy bands influences the source identification capabilities. In particular, the fraction of detected neutrino events above a given energy threshold plays a key role in the achievable signal-to-noise ratio: astrophysical neutrinos from AGN are expected to dominate at higher energies, while the atmospheric background falls steeply. Moreover, higher-energy neutrinos are reconstructed with better angular resolution, further enhancing source identification capabilities.
@@
Figure 3 illustrates the trade-off between neutrino energy threshold and exposure time for future neutrino observatories, with the left panel corresponding to IceCube-Gen2 and the right panel to the proposed HUNT detector. The horizontal axis denotes the minimum reconstructed neutrino energy threshold applied to the event sample, while the vertical axis indicates the observation time required to reach a given detection significance. For IceCube-Gen2, the energy-dependent angular resolution is adopted from the projected point-spread function reported in ice2_psf2019, while for the HUNT detector the calculation incorporates its energy-dependent angular resolution and effective area (private communication with Tian-Qi Huang).
@@
As the energy threshold of neutrinos increases, two competing effects emerge: (i) the angular resolution improves and the atmospheric neutrino flux drops; (ii) the number of neutrino events above the energy threshold declines, reducing statistical power. These opposing trends jointly define an optimal energy window, typically around a few hundred TeV, where the sensitivity peaks. Beyond this point, further increases in the energy threshold yield diminishing returns due to low event rates. Owing to the expected angular resolution of the HUNT detector being improved by a factor of $\sim$4–5 relative to IceCube-Gen2, together with an effective area larger by a factor of $\sim$3, the HUNT detector naturally achieves a higher detection significance for a given energy threshold and exposure time.
V Impact of Systematic and Positional Uncertainties
@@
The analyses presented thus far assumed idealized conditions: the AGN X-ray luminosity function (XLF) was treated as perfectly known, and the neutrino angular uncertainty was assumed to be fixed for all events. In practice, both assumptions introduce additional uncertainties in the inferred signal fraction $f$. Here we quantify the contributions from these two dominant sources of uncertainty and evaluate their combined impact on the overall detection significance.
@@
The first source of uncertainty arises from the extrapolation of the bright AGN subsample (above $F_{\rm min}$) to the entire AGN population. In our baseline approach, this extrapolation is carried out using the luminosity distribution of the eFEDS AGN catalog, which provides direct observational data and avoids the need for an explicit parametric model. However, the eFEDS sample does not fully capture the global AGN population, particularly at low luminosities and high redshifts. To assess this systematic, we perform an alternative extrapolation based on the parametric AGN XLF. Assuming a linear relation between neutrino and X-ray luminosities (e.g., Abbasi2022PhRvD; Kun2024PhRvD), the total neutrino flux is expressed as
@@
where $\phi(L_{X},z)$ is the XLF. Using the best-fit parameters and reported uncertainties from recent XLF measurements (UedaXLF2014), we propagate parameter variations through Monte Carlo simulations to estimate the relative uncertainty in the extrapolated neutrino contribution as a function of the X-ray flux threshold $F_{\rm min}$. The resulting uncertainty is summarized in Figure 4. As expected, adopting brighter flux thresholds leads to significantly larger uncertainties. This arises because selecting only the most luminous AGNs (e.g., $F_{\rm min}>3\times 10^{-13}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$) leaves the majority of the AGN population unresolved; extrapolation from such a small subset amplifies the impact of uncertainties in the faint-end slope and normalization of the XLF, potentially causing the correction factor to vary by factors of several. In these extreme cases, the extrapolation becomes unreliable and the inferred total neutrino contribution cannot be robustly constrained. Therefore, this method is only recommended for flux thresholds below $\sim 2\times 10^{-13}\ \mathrm{erg\ cm^{-2}\ s^{-1}}$, where the extrapolation remains well-behaved.
@@
Figure 4: Relative uncertainty factor in the extrapolated AGN contribution as a function of the flux threshold $F_{\rm min}$. The blue line shows the mean correction factor, and shaded regions indicate $1\sigma$, $2\sigma$, and $3\sigma$ confidence intervals. Uncertainties remain modest at low thresholds but grow rapidly when only the brightest AGNs are included.
@@
The second source of uncertainty stems from event-by-event variations in the $\sigma_{\rm psf}$. In previous sections, we adopted a fixed PSF width $\sigma_{\rm psf}$ for a given energy bin, whereas real detectors exhibit a distribution of reconstructed angular uncertainties. This variation effectively broadens the signal and background probability density functions used in the maximum-likelihood analysis, thereby reducing sensitivity. To quantify this effect, we simulate neutrino samples by assigning to each event a $\sigma_{\rm psf}$ randomly sampled from the detector’s energy-dependent angular resolution distribution.
Combining the uncertainties from both XLF extrapolation and PSF variation, we assess the sensitivity of future neutrino telescopes to testing the AGN-origin hypothesis. As a more generic and near-term scenario, we first consider a 10 km3 undersea neutrino telescope with an angular resolution of $0.2^{\circ}$, representative of IceCube-Gen2. We include all neutrino events with energies above 100 TeV and select the full AGN sample with X-ray fluxes $F_{X}>2.4\times 10^{-13}\,\mathrm{erg\,s^{-1}\,cm^{-2}}$. The uncertainty associated with the extrapolation of the XLF is modeled as a relative error of 30%. Under these assumptions, such a detector would achieve a significance of approximately $2.4\sigma$ after five years of exposure in testing whether all astrophysical neutrinos originate from AGNs, as shown in the Figure 5. This level of significance, while suggestive, is insufficient for a decisive test of the AGN-origin hypothesis.
Figure 5: Expected Detection significance as a function of exposure time for different detector configurations and energy selections. The vertical dotted line indicates an exposure of 5 yr, with the corresponding $(S/N)_{f}$ values marked for comparison.
As an illustrative and more optimistic case, we consider a 30 km3 neutrino detector and focus on events with energies above 300 TeV, where an angular resolution of $0.04^{\circ}$ is expected. At this energy threshold, we restrict the analysis to AGNs with X-ray fluxes $F_{X}>4\times 10^{-14}\,\mathrm{erg\,s^{-1}\,cm^{-2}}$. The uncertainty arising from the extrapolation of the X-ray luminosity function is again treated as a relative error of 10%, and variations in the detector PSF are also taken into account. Under this configuration, a 30km3 detector can test the AGN-origin hypothesis at a confidence level of roughly $8.6\sigma$ after three years of exposure, even after accounting for both XLF extrapolation and $\sigma_{\rm psf}$ variation uncertainties, as shown in the Figure 5. We further explore a broader energy range of above 100 TeV. In this case, the expected significance is $8\sigma$ after five years of exposure. The limited improvement despite increased statistics highlights the dominant role of XLF extrapolation uncertainties, which ultimately constrain further sensitivity gains.
VI Discussion
Figure 6: Left: Signal-to-noise ratio $(S/N)_{f}$ as a function of neutrino fraction $f$, showing that even partial AGN contributions can be constrained with high significance. Right: $(S/N)_{f}$ as a function of the flux-weighting parameter $s$ ($N_{\nu}\propto F_{X}^{s}$), illustrating that stronger weighting toward luminous AGNs enhances detectability. All panels correspond to a 30 km3 undersea neutrino telescope with 3 years of exposure.
VI.1 AGN as a Fractional Source of the Diffuse neutrino Flux
@@
Throughout most of this work we have illustrated the methodology using the limiting case in which AGNs are assumed to be the dominant contributors to the observed astrophysical neutrino flux. In reality, the diffuse neutrino background could be produced by a mixture of source populations, and AGNs may account for only a fraction of the total. Our framework readily accommodates this more general scenario: the parameter $f$ in the unbinned maximum-likelihood formalism already represents the fraction of events associated with AGNs, while the remaining (1f)(1-f) fraction is treated as isotropic background.
Figure 6 (left) demonstrates how the signal-to-noise ratio ${(S/N)}_{f}$ scales with $f$ for a representative next-generation detector scenario (a 30 km3-scale underwater neutrino telescope with 3 years of exposure). The trend is nearly linear, indicating that even modest AGN contributions (e.g., $f\sim 0.2$) remain detectable at $\sim 4\sigma$ significance. This scaling behavior arises because the likelihood ratio effectively isolates the spatially correlated AGN component, which diminishes proportionally with $f$, while the statistical background remains unchanged.
VI.2 X-ray flux-weighting

### v3

@@ III Predictive Power for Future Neutrino Detectors
Figure 3 illustrates how the detection significance depends on the angular resolution and the total number of detected neutrinos, where the astrophysical neutrino fraction above 100 TeV is fixed at $f=0.5$, consistent with the IceCube measurements. (Abbasi et al., 2022b) The trends are clear: improving the angular resolution reduces positional uncertainties, thereby increasing the contrast between AGN–neutrino associations and random coincidences. At a fixed angular resolution, increasing the total event count enhances the statistical weight of the likelihood, roughly following the expected ${(S/N)}_{f}\propto\sqrt{N}$ scaling. On the other hand, for a fixed number of detected neutrinos, the significance decreases with increasing PSF width by ${(S/N)}_{f}\propto\sigma_{\rm psf}^{-1}$. Notably, achieving a $5\sigma$ detection requires only $\sim 2000$ events for an angular resolution of $0.1^{\circ}$, whereas resolution worse than $0.4^{\circ}$ demands at least two orders of magnitude more events, making identification of the correlation unattainable. As marked with the red star in Figure 3, with $\sim 10^{2}$$10^{3}$ detected astrophysical neutrinos and a median angular uncertainty of $\sim 0.55^{\circ}$ above 100 TeV(Zegarelli et al., 2025), IceCube’s observations do not have sufficient exposure and angular resolution to diagnosing the AGN origin of diffuse TeV neutrinos: the current IceCube data lies far below the regime where AGN population correlations could be statistically established (denoted by ”IceCube” in Figure 3). This highlights the importance of angular resolution. This trend can be understood as the combined result of two complementary effects acting at different levels of the analysis. First, at the population level, a smaller $\sigma_{\rm psf}$ mitigates source confusion and allows the likelihood analysis to incorporate contributions from a larger fraction of the AGN population, including sources with weaker X-ray emission, provided that the mean source separation satisfies $\bar{d}\geq 3\sigma_{\rm psf}$. In this regime, reducing $\sigma_{\rm psf}$ allows more sources to satisfy $\bar{d}\geq 3\sigma_{\rm psf}$, so that they can be included in the likelihood with non-negligible weights (the $w_{j}$ of Eq. 5), thereby enhancing the cumulative statistical power. Second, at the event level, improved angular resolution sharpens the spatial probability density of each neutrino event, concentrating the signal likelihood around the true source positions. This reduces the overlap with the isotropic background term and increases the contrast between genuine AGN–neutrino associations and random coincidences (i.e., the $S_{i}$ and $B_{i}$ of Eq. 4), leading to a larger likelihood ratio and a higher detection significance.
@@
Figure 4: Trade-off between neutrino energy threshold and exposure time for future neutrino observatories. The left panel shows results for IceCube-Gen2, adopting the projected energy-dependent angular resolution from Bradascio and Glüsenkamp (2019). The right panel shows the corresponding results for the proposed 30 km3 HUNT detector, incorporating its energy-dependent angular resolution and effective area from Huang and HUNT Collaboration (2025).
@@
Building upon the angular resolution-based considerations described above, we next examine how the selection of different neutrino energy bands influences the source identification capabilities. In particular, the fraction of detected neutrino events above a given energy threshold plays a key role in the achievable detection significance: astrophysical neutrinos from AGN are expected to dominate at higher energies, while the atmospheric background falls steeply. Moreover, higher-energy neutrinos are reconstructed with better angular resolution, further enhancing source identification capabilities.
@@
Figure 4 illustrates the trade-off between neutrino energy threshold and exposure time for future neutrino observatories, with the left panel corresponding to IceCube-Gen2 and the right panel to the proposed HUNT detector. The horizontal axis denotes the minimum reconstructed neutrino energy threshold applied to the event sample, while the vertical axis indicates the observation time. For IceCube-Gen2, we adopt an energy-dependent angular resolution based on muon track reconstruction reported in Bradascio and Glüsenkamp (2019), together with the corresponding effective area taken from the official IceCube-Gen2 design study (Aartsen et al., 2021). For the HUNT detector, the calculation incorporates its energy-dependent angular resolution and effective area from the simulations (Huang and HUNT Collaboration, 2025).
@@
As the energy threshold of neutrinos increases, two competing effects emerge: (i) the angular resolution improves and the atmospheric neutrino flux drops; (ii) the number of neutrino events above the energy threshold declines, reducing statistical power. These opposing trends jointly define an optimal energy window, typically around a few hundred TeV, where the sensitivity peaks (100 TeV for IceCube-Gen2 and 300 TeV for HUNT). Beyond this regime, further increases in the energy threshold result in diminishing returns due to limited event statistics.
Within the same physical framework, the superior performance of the HUNT detector can be understood in terms of its expected design characteristics. Based on its projected, energy-dependent performance, HUNT is anticipated to achieve a median angular resolution that is improved by a factor of $\sim$4–5 relative to IceCube-Gen2, together with an effective area larger by a factor of $\sim$3 in the relevant energy range (a few hundred TeV). These factors summarize the relative detector performance inferred from their respective design specifications, rather than representing a direct rescaling of IceCube-Gen2 parameters. As a result, HUNT attains a higher detection significance—or equivalently, requires a shorter exposure time—for a given neutrino energy threshold.
IV Impact of Systematic and Positional Uncertainties
@@
The analyses presented thus far have assumed idealized conditions: the AGN X-ray luminosity function (XLF) was treated as perfectly known, and the neutrino angular uncertainty was assumed to be fixed for all events. In practice, both of them have uncertainties, so they could introduce additional uncertainties in the inferred signal fraction $f_{\rm tot}$. Here we quantify the contributions from these two dominant sources of uncertainty and evaluate their combined impact on the overall detection significance. In addition to these systematic effects, statistical fluctuations due to the finite neutrino sample are automatically incorporated in all Monte Carlo realizations, and are not treated as a separate source of uncertainty. As the exposure increases, the impact of these statistical fluctuations diminishes, and the total uncertainty becomes increasingly dominated by systematic effects.
@@
The first source of the systematic uncertainty arises from extrapolating the neutrino contribution inferred from the bright, flux-limited AGN subset ($F_{X}>F_{\rm min}$) to the full AGN population. In our baseline analysis, this extrapolation is implemented using the observed X-ray flux distribution of the eFEDS AGN catalog, which provides a data-driven correction without assuming a specific parametric form. However, the eFEDS sample does not fully represent the global AGN population, particularly toward low luminosities and high redshifts, where the bulk of the luminosity density resides.
To quantify the systematic uncertainty associated with this extrapolation, we alternatively model the AGN population using a parametric XLF. Assuming that the neutrino luminosity scales linearly with the X-ray luminosity (e.g., Abbasi et al., 2022a; Kun et al., 2024), the total neutrino flux can be written as
@@
where $\phi(L_{X},z)$ denotes the XLF.
Using the best-fit XLF parameters and their reported $1\sigma$ uncertainties from Ueda et al. (2014), we perform Monte Carlo simulations in which the XLF parameters are randomly varied according to independent Gaussian distributions centered on their best-fit values. For each realization, we compute the corresponding correction factor between the flux-limited subset and the full population, thereby estimating the relative uncertainty of the extrapolated neutrino contribution as a function of the adopted flux threshold $F_{\rm min}$. That quantifies the relative uncertainty in the extrapolated total AGN neutrino contribution induced solely by XLF uncertainties, which we denote as $\delta_{\rm XLF}$. The resulting $\delta_{\rm XLF}$ as a function of $F_{\rm min}$ is summarized in Figure 5. As expected, adopting brighter flux thresholds leads to rapidly increasing extrapolation uncertainties. This behavior reflects the fact that when only very luminous AGNs are included, the selected subset captures only a small fraction of the total XLF-weighted luminosity density, causing the correction factor to become highly sensitive to uncertainties in the faint-end slope and normalization of the XLF.
As discussed in Sec. II.2, the total AGN neutrino fraction $f_{\rm tot}$ is obtained by applying a flux-based extrapolation to the flux-limited likelihood result, as defined in Eq. (3). Uncertainties in the XLF-based extrapolation propagate directly into the inferred total AGN neutrino fraction $f_{\rm tot}$. In the Monte Carlo framework, this appears as a broadening of the reconstructed $f_{\rm tot}$ distribution, characterized by an increased standard deviation $\sigma_{f_{\rm tot}}$, while the mean remains approximately unbiased. We quantify the impact of this effect using the relative uncertainty $\delta$. When $\delta>50\%$, the inference of $f_{\rm tot}$ becomes dominated by XLF systematic uncertainty rather than statistical fluctuations, and the extrapolated AGN contribution is no longer robustly constrained. To avoid this regime, we restrict our analysis to flux thresholds $F_{\rm min}<7\times 10^{-13}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$, which corresponds to the stability criterion quantified above, as shown in Figure 5. In this range, the selected AGN subset encompasses a substantial fraction of the total XLF-weighted luminosity density, and the extrapolation remains stable against variations in the XLF parameters.
@@
Figure 5: Relative uncertainty factor $\delta_{\rm XLF}$ of XLF in the extrapolated AGN contribution as a function of the AGN flux threshold $F_{\rm min}$. The shaded band shows the 68% confidence interval derived from Monte Carlo realizations. The vertical blue, orange, and green lines mark the values of $F_{\rm min}$ adopted for IceCube-Gen2 ($E_{\nu}>100$ TeV), HUNT ($E_{\nu}>100$ TeV), and HUNT ($E_{\nu}>300$ TeV), respectively. Uncertainties remain modest at low thresholds but grow rapidly when only the brightest AGNs are included.
@@
The second uncertainty arises from the event-by-event variations in the $\sigma_{\rm psf}$. In previous sections, we adopted a fixed PSF width $\sigma_{\rm psf}$ for given energy bins, whereas real detectors exhibit a distribution of reconstructed angular uncertainties. This variation effectively broadens the signal and background probability density functions used in the maximum-likelihood analysis, thereby reducing sensitivity. To quantify this effect, we simulate neutrino samples by assigning to each event a $\sigma_{\rm psf}$ randomly sampled from the detector’s energy-dependent angular resolution distribution. We find that the uncertainty induced by event-by-event PSF variations is smaller than that from the XLF-based extrapolation, and therefore does not dominate the total error budget.
Combining the uncertainties from both XLF extrapolation and PSF variation, we assess the sensitivity of future neutrino telescopes to testing the AGN-origin hypothesis. In the simulations, the angular uncertainty of each neutrino event is treated on an event-by-event basis, with the reconstructed direction randomly drawn according to the corresponding point-spread function, thereby accounting for realistic variations in the localization accuracy. Focusing on events with energies above 100 TeV ($\beta\simeq 0.5$), IceCube-Gen2 is expected to achieve a median angular resolution of $\sim 0.2^{\circ}$. To avoid source confusion, we therefore require the selected AGN population to satisfy the condition $\bar{d}\geq 0.6^{\circ}$. Based on the eFEDS AGN catalog over the $142\,\mathrm{deg}^{2}$ survey area, this requirement translates into a maximum of $\sim 150$ AGNs within the field of view. This criterion selects the brightest sources with X-ray fluxes $F_{X}>2.4\times 10^{-13}\,\mathrm{erg\,s^{-1}\,cm^{-2}}$, which are sufficiently sparse to remain well resolved at the IceCube-Gen2 angular resolution. As shown in Figure 5, for the flux threshold corresponding to IceCube-Gen2 (blue vertical line), the uncertainty induced by the XLF-based extrapolation from the flux-limited AGN subsample to the full AGN population is $\delta_{\rm XLF}\simeq 30\%$. As shown in Figure 6, such the detector would achieve a significance of approximately $\sim 3.1\sigma$ after five years of exposure in testing whether all astrophysical neutrinos originate from AGN population. This significance is suggestive but insufficient for a decisive test of the AGN-origin hypothesis.
Figure 6: Top panel: Relative uncertainty on the inferred total AGN neutrino fraction $\delta$ as a function of the exposure time, derived from repeated Monte Carlo simulations. Bottom panel: The detection significance, expressed as $\sqrt{\rm TS}$, as a function of exposure time. The steady decrease of $\delta$ and the growth of $\sqrt{\rm TS}$ demonstrate how both the precision and statistical significance of the AGN contribution improve with increased exposure.
As an illustrative and more optimistic case, we consider the HUNT, a proposed 30 km3 underwater neutrino detector (Huang et al., 2024; Huang and HUNT Collaboration, 2025) and focus on events with energies above 300 TeV ($\beta\simeq 0.8$), for which an angular resolution of $\sim 0.04^{\circ}$ is expected at these energies. At this energy threshold, we restrict the analysis to AGNs with X-ray fluxes $F_{X}>4\times 10^{-14}\,\mathrm{erg\,s^{-1}\,cm^{-2}}$. Based on the eFEDS AGN catalog, this $\bar{d}\geq 3\sigma_{\rm psf}$ condition corresponds to a maximum surface density of $\sim 3200$ sources, which in turn selects the brightest AGNs with $F_{X}>4\times 10^{-14}\,\mathrm{erg\,s^{-1}\,cm^{-2}}$. Such a relatively low flux threshold significantly reduces the uncertainty associated with the XLF-based extrapolation. As indicated by the green vertical line in Figure 5, the corresponding relative uncertainty is $\delta_{\rm XLF}\simeq 10\%$. We find that HUNT can test the AGN-origin hypothesis at a confidence level of roughly $8.7\sigma$ after one years of exposure, after accounting for uncertainties from both XLF extrapolation and PSF variation, as shown in Figure 6. At the same time, the likelihood analysis constrains the total AGN neutrino fraction to a relative precision of $\delta\simeq 20\%$. With five years of exposure, the significance further increases to $\sim 19\sigma$, and the relative uncertainty decreases to $\delta\simeq 10\%$.
For comparison, we also consider a more conservative energy threshold of $E_{\nu}>100$ TeV for HUNT. In this case, the larger angular uncertainty requires a brighter AGN flux threshold (orange vertical line in Figure 5), leading to a larger XLF extrapolation uncertainty and consequently a lower detection significance (orange curve in Figure 6), although the AGN-origin hypothesis remains testable at high confidence with multi-year exposure.
V Discussion
V.1 AGNs contributing partially to the total astrophysical neutrino flux
Figure 7: Top panel: Recovered AGN neutrino fraction $f_{\rm tot}$ as a function of the assumed AGN neutrino fraction $\alpha$. The blue shaded band shows the central 68% interval from 800 Monte Carlo realizations, while the solid blue line indicates the mean of $f_{\rm tot}$ from the simulations. The red dashed line represents the theoretical expectation $f_{\rm tot}=\alpha\times\beta$, showing good agreement between the likelihood reconstruction and the expected value. Bottom panel: Corresponding detection significance $\sqrt{\rm TS}$ as a function of $f_{\rm tot}$, showing that even partial AGN contributions can be detected with high statistical significance.
@@ V.1 AGNs contributing partially to the total astrophysical neutrino flux
Throughout most of this work we have assumed the limiting case in which AGNs are assumed to be the dominant contributors to the observed astrophysical neutrino flux. In reality, the diffuse neutrinos could be produced by a mixture of source populations, and AGNs may account for only a fraction of the total astrophysical neutrino flux. Our framework readily accommodates this more general scenario: the parameter $f_{\rm tot}$ in the unbinned maximum-likelihood formalism already represents the fraction of events associated with AGNs, while the remaining $(1-f_{\rm tot})$ fraction is treated as isotropic background.
[…]
V.2 X-ray flux-weighting
Figure 8: Top panel: Relative uncertainty of the recovered AGN neutrino fraction $\delta$ as a function of the flux-weighting parameter $s$ ($N_{\nu}\propto F_{X}^{s}$). Bottom panel: Detection significance $\sqrt{\rm TS}$ as a function of the flux-weighting parameter $s$, illustrating that stronger weighting toward luminous AGNs enhances the detectability. All panels correspond to a 30 km3 undersea neutrino telescope with 3 years of exposure.